W. Edwards Deming in his red bead experiment had a box of 4,000 beads, of which 800 were red and 3,200 were white. Suppose a researcher were to conduct a modified version of the red bead experiment. In her experiment, she has a bag of 20 beads, of which 4 are red and 16 are white. This experiment requires a participant to reach into the bag and randomly select five beads, one at a time, without replacement.?

What is the probability that the participant will select exactly four white beads?What is the probability that the participant will select exactly four red beads? What is the probability that the participant will select all red beads??
Answer #1

A) P(selecting 4 white beads) = 16C4 * 4C1/20C5 = 0.47

b) P(exactly 4 red beads) = 4C4 * 16C1/20C5 = 0.001

c) P(selecting all red beads) = 4C4 * 16C1/20C5 = 0.001